OpenStudy (anonymous):

write an equation that relates x and y. use your equation to find the value of y when x=15.

OpenStudy (anonymous):

y=mx+b f(x) = 5x + 2 f(15) = 5(15) +2

OpenStudy (anonymous):

thank you!